# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 9.2. Strict comparison with minimal monotonicity and quasi-convex functions
    Consider now a vanishing sequence $\varepsilon_{k}\searrow 0$. By the compactness of $\overline{\Omega}$, up to a subsequence, $y_{k}\vcentcolon=y_{\varepsilon_{k}}\to\overline{y}\in\partial\Omega$ as $k\to\infty$. Set $\delta_{k}:=\delta(\varepsilon_{k})$ and note that, by (9.21), one has
@@
    Now, since $\delta_{k}\searrow 0$ and $y_{k}\to\overline{y}\in\partial\Omega$ as $k\to+\infty$, for each $\rho>0$ fixed one has $B_{\delta_{k}}(y_{k})\subset B_{\rho}(\overline{y})$ if $k$ is large enough. Hence, taking the limsup in (LABEL:sc_bound), one obtains
    $\limsup_{k\to\infty}u_{m}^{\varepsilon_{k}}(y_{k})\leqslant\max_{\overline{B}_{\rho}(\overline{y})}u_{m}\qquad\forall\rho>0$
@@
    10. From the Euclidean space to manifolds: a brief note
    In our presentation we decided to remain in the Euclidean context. However, many of the results discussed here extend to Riemannian manifolds; for example, the Almost Everywhere Theorem 8.1, the Subharmonic Addition Theorem 8.3 and the Strict Comparison Theorem 9.17 have been given in [26] in a more general form on manifolds. We note that, roughly speaking, all local results naturally extend from Euclidean spaces to manifolds, via a suitable use of local coordinates.
    In this passage, one needs to extend the definition of quasi-convex function to real-valued functions defined on a manifold $M$. The usual definition of a convex function on a Riemannian manifold is that of a function which is convex along each geodesic; that is, $f\colon M\to{\mathbb{R}}$ is convex on a geodesically convex subset $U\subset M$ if and only if $f\circ\gamma\colon(-\varepsilon,\varepsilon)\to{\mathbb{R}}$ is convex (in the real sense) for every geodesic segment $\gamma\colon(-\varepsilon,\varepsilon)\to U$. Hence, by paraphrasing the idea that a quasi-convex function is convex modulo a quadratic perturbation, one can give, on connected manifolds, the following definition:
    $f$ is $2\lambda$-quasi-convex if $f+\lambda\,d(\cdot,p)^{2}$ is convex for all $p\in M$,
    where $d$ is the standard metric on $M$ induced by the Riemmann structure.
    When $M={\mathbb{R}}^{n}$ is a Euclidian space, this definition is equivalent to the one we have given in ${\mathbb{R}}^{n}$. Alternatively, one could only ask that $f+\lambda\;\!d(\cdot,p)^{2}$ be convex for at least one point $p\in M$. The two definitions are clearly equivalent in the Euclidean case, and in general on every manifold with zero sectional curvature. However, the first definition is, in general, stronger and it is the right definition for many purposes. For example, the stronger definition gives the extension to Riemannian manifolds of the characterization
@@ 10. From the Euclidean space to manifolds: a brief note
    as noted in [3]. [^12] Let us also note that on manifolds with positive curvature it could be tedious that $d(\cdot,p)^{2}$ is not convex; for instance, one cannot say that $f$ being $\lambda$-quasi-convex implies that $f$ is $\mu$-quasi-convex for all $\mu>\lambda$.
    On the other hand, often one needs only local quasi-convexity, a notion which requires some additional explanation. One makes use of the normal neighborhood theorem (see, e.g., Theorem 9.14 of [44]) which states that for every $p_{0}\in M$ and $\varepsilon>0$ there exists a neighborhood $W$ of $p_{0}$ such that each pair of points $q,q^{\prime}\in W$ is connected by a unique minimal geodesic parametrized on $[0,1]$ and of length less than $\varepsilon$; namely, $\gamma_{q,q^{\prime}}(t)=\exp_{q}(tv)$ for the unique $v=v(q,q^{\prime})\in T_{q}M$ with $|v|<\varepsilon$. For such a normal neighborhood $W$, one then can define
@@
    This leads to a natural extension of the definition of local quasi-convexity. Notice that, in a normal chart $(U,\varphi)$ about $q\in M$, [^13] That is, a coordinate chart $(U,\varphi)$ with $U=\exp_{q}(B_{\varepsilon}(0))$ and $\varphi=\psi\circ\exp^{-1}$, for some linear isometry $\psi\colon T_{q}M\to{\mathbb{R}}^{m}$, where $m=\mathrm{dim}\,M$. the function $f\circ\varphi^{-1}$ is $2\lambda$-quasi-convex along each small linear segment starting from $0\in{\mathbb{R}}^{m}$. This partially motivates why one can also rightfully say, without relying on the metric, that
@@
    Adopting this definition of local quasi-convexity, one straightforwardly extends to Riemannian manifolds the results mentioned at the beginning of the section.
    Acknowledgements
    The authors express their deep gratitude to Blaine Lawson and Reese Harvey for their generosity and encouragement to perform and updated and enriched synthesis of the unpublished manuscripts [25] and [26]. In addition, numerous useful conversations with Reese Harvey have contributed to the development of this work. The authors are partially supported by the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM), Italy. Redaelli partially supported by Fondazione Cassa di Risparmio di Padova e Rovigo, Italy.
    Appendix A.
    \@afterheading
@@ Acknowledgements
    Lt:aboveremark Notice that the definition (LABEL:leg:def2) is somewhat implicit; that is, in order to define $g(y)=\mathscr{L}f(y)$ for a given $y\in\partial f(X)={\rm dom}(g)$, one must identify a fiber $\partial f(x)$ in which $y$ lives. A more explicit way to define the Legendre transform $g$ of $f$ involves an optimization procedure based on the observation that
    $(x,y)\in\partial f\quad\iff\quad f(x)+\langle y,z-x\rangle\leqslant f(z),\quad\forall z\in X,$
    or equivalently
    $f(z)-\langle y,z\rangle\geqslant f(x)-\langle y,x\rangle,\quad\forall z\in X.$
@@
    We will see that these equivalent formulations (LABEL:leg:def2) and (A.3) are both useful.
@@
    Let $f:X\to{\mathbb{R}}$ be convex on $X\subset{\mathbb{R}}^{n}$ open and convex. Then the following properties hold.
@@
      $\mathscr{L}f(y)=\sup_{z\in X}\left(\langle y,z\rangle-f(z)\right);$ (A.4)
@@
      for each $x\in X$, $y\in\partial f(X)$,
      $(x,y)\in\partial f\ \ \iff\ \ (y,x)\in\partial\mathscr{L}f;$
@@
      $\mathscr{L}$ is an involution; that is, $\mathscr{L}(\mathscr{L}f)\big{|}_{X}=f$.
@@
    The equivalence of part (a) has been presented in LABEL:Lt:aboveremark.
    The convexity of $\mathscr{L}f$ in part (b) now follows from the fact that $\mathscr{L}f$ is the supremum of the family of affine functions $\{\langle\cdot,z\rangle-f(z)\}_{z\in X}$.
    For the implication ()($\Longrightarrow$)( ⟹ ) of part (c), let $(x,y)\in X\times\partial f(X)$. Suppose that $(x,y)\in\partial f$ (that is, $y\in\partial f(x)$). In order to show that $(y,x)\in\partial g$ where $g=\mathscr{L}f$, one needs to show that $y\in\partial f(x)$; that is,
@@
    $f(z)+\langle w,\hat{x}-z\rangle\leqslant f(\hat{x}),\quad\forall\,\hat{x}\in X.$ (A.6)
    By the definition (LABEL:leg:def2) of the Legendre transform we have
    $w\in\partial f(z)\ \ \Rightarrow\ \ g(w):=\langle z,w\rangle-f(z)$
    and
    $y\in\partial f(x)\ \ \Rightarrow\ \ g(y):=\langle x,y\rangle-f(x)$
    which imply
@@
    where the last inequality uses (A.6) in $\hat{x}=x$ and gives (A.5). Notice that we also have $X\subset\partial g(\partial f(X))$ since $x\in X$ is arbitrary and $y\in\partial f(x)\neq\emptyset$ implies $x\in\partial g(y)$.
    For the involution claim of part (d), notice that we have shown that $(x,y)\in\partial f$ implies that $(y,x)\in\partial\mathscr{L}f=\partial g$, where $g$ is convex. Hence, for each $x\in X$, by the definition (LABEL:leg:def), we have
    $\mathscr{L}(\mathscr{L}f)(x)=\mathscr{L}g(x):=\langle x,y\rangle-g(y):=\langle x,y\rangle-\langle y,x\rangle+f(x)=f(x),$
    which proves (d) since $x\in X$ is arbitrary.
@@
    $\langle x,y\rangle=\mathscr{L}f(y)+\mathscr{L}\!\mathscr{L}f(y)=\mathscr{L}f(y)+f(x)$
@@
    $\langle x,y\rangle=\sup_{z\in X}\left(\langle y,z\rangle-f(z)\right)+f(x)\geqslant\langle y,z\rangle-f(z)+f(x),\quad\forall z\in X,$
    yielding $f(z)\geqslant f(x)+\langle y,z-x\rangle$ for each $z\in X$; that is, $y\in\partial f(x)$, which is equivalent to $(x,y)\in\partial f$. ∎
@@
    It is false in general that
    $(x,y)\in\partial f\ \iff\ (y,x)\in\partial g.$
@@
    $\partial f(x)=\begin{cases}-1&\text{if $x\in(-1,0)$}\\ [-1,1]&\text{if $x=0$}\\ 1&\text{if $x\in(0,1)$}\end{cases},$
    so that, according to (LABEL:leg:def)
    $g\vcentcolon=\mathscr{L}f\equiv 0\quad\text{on}\quad\partial f(X)=[-1,1].$
    Therefore,

### v3

@@ 9.3. Strict comparison with minimal monotonicity and semiconvex functions
Consider now a vanishing sequence $\varepsilon_{k}\searrow 0$. By the compactness of $\overline{\Omega}$, up to a subsequence, $y_{k}\vcentcolon=y_{\varepsilon_{k}}\to\overline{y}\in\partial\Omega$ as $k\to\infty$. Set $\delta_{k}\vcentcolon=\delta(\varepsilon_{k})$ and note that, by (9.16), one has
@@
Now, since $\delta_{k}\searrow 0$ and $y_{k}\to\overline{y}$, for each $\rho>0$ fixed one has $B_{\delta_{k}}(y_{k})\subset B_{\rho}(\overline{y})$ if $k$ is large enough. Hence, taking the limsup in (9.17), one obtains
$\limsup_{k\to\infty}u_{m}^{\varepsilon_{k}}(y_{k})\leqslant\max_{\overline{B}_{\rho}(\overline{y})}u_{m}\qquad\forall\rho>0,$
@@
Chapter 10 From Euclidean spaces to manifolds: a brief note
In this chapter we show how the notion of semiconvexity can be transported from Euclidian spaces to manifolds. The key point will be to show that the notion of semiconvexity is preserved under smooth coordinate changes.
10.1. Objectives
In our presentation, for simplicity, we have worked only in Euclidean spaces. However, many of the results discussed here extend to manifolds; for example, the Almost Everywhere Theorem 8.1, the Subharmonic Addition Theorem 8.3 and the Strict Comparison Theorem 9.19 have been given in [29] in a more general form on manifolds. We note that, roughly speaking, all local results naturally extend from Euclidean spaces to manifolds, via a suitable use of local coordinates.
To make this passage, one needs to extend the definition of a locally semiconvex function on an open subset of ${\mathbb{R}}^{n}$ (see Definition 2.1) to functions defined on a manifold $M$ of dimension $n$. The crucial observation for this extension is to prove that the class of locally semiconvex functions on an open subsets of ${\mathbb{R}}^{n}$ is stable under pull-backs via diffeomorphisms (that is, smooth changes of coordinates).
We will prove this claim and then formulate a definition of a locally convex function on a manifold.
10.2. Invariance of local semiconvexity under coorinate changes
The discussion above is formalized in the following statement.
Theorem 10.1.
Suppose that $\Phi:X\to Y$ is a smooth diffeomorphism between open subsets $X,Y\subset{\mathbb{R}}^{n}$. Then $\Phi$ pulls back locally semiconvex functions on $Y$ to locally semiconvex functions on $X$; that is,
$u:Y\to{\mathbb{R}}$ is locally semiconvex  $\implies\quad u\circ\Phi:X\to{\mathbb{R}}$ is locally semiconvex.
We are going to reduce Theorem 10.1 to Lemma 10.3 below. This reduction involves several simplifications, where the final step will be a potential-theoretic argument which exploits the decreasing limit property and smooth approximations (mollifications) that preserve convexity and Lipschitz constants. The properties of the mollifications that we will need are contained in the following result which augments Lemma 2.4. We make use of an even mollifier; that is,
@@ 10.2. Invariance of local semiconvexity under coorinate changes
and
$\eta(-z)=\eta(z)\quad\forall z\in{\mathbb{R}}^{n}.$ (10.1)
Lemma 10.2.
Suppose that $u$ is convex and $K$-Lipschitz on a convex neighborhood $\mathcal{N}(\overline{Y})$ of $\overline{Y}$ convex. For each $\varepsilon>0$, consider the mollified function
$u_{\varepsilon}(y)\vcentcolon=(u\ast\eta_{\varepsilon})(y)=\int_{{\mathbb{R}}^{n}}u(y-z)\eta_{\varepsilon}(z)\,\mathrm{d}z=\int_{{\mathbb{R}}^{n}}u(y-z)\eta(z/\varepsilon)\,\varepsilon^{-n}\,\mathrm{d}z,$ (10.2)
for $y\in Y^{\varepsilon}\vcentcolon=\{y\in\mathcal{N}(\overline{Y}):\ {\rm dist}(y,\partial\mathcal{N}(\overline{Y}))>\varepsilon\}$ is convex and contains $\overline{Y}$ for all sufficiently small $\varepsilon$. The following hold:
  • (a)
    $u_{\varepsilon}$ is smooth and convex on $Y^{\varepsilon}$;
  • (b)
    $u_{\varepsilon}$ is $K$-Lipschitz on $Y^{\varepsilon}$;
  • (c)
    $u_{\varepsilon}\searrow u$ on compact subsets of $\mathcal{N}(\overline{Y})$ as $\varepsilon\searrow 0$
Proof.
By the condition on the support of $\eta$, $u_{\varepsilon}$ defined by (10.2) can also be expressed as
$u_{\varepsilon}(y)\vcentcolon=\int_{B_{\varepsilon}(0)}u(y-z)\eta_{\varepsilon}(z)\,\mathrm{d}z$ (10.3)
and hence is clearly well-defined on $Y^{\varepsilon}$, which contains $\overline{Y}$ for all small $\varepsilon$. The smoothness of $u_{\varepsilon}$ on $Y^{\varepsilon}$ is a standard consequence of the dominated convergence theorem for $u$ continuous. The convexity of $u_{\varepsilon}$ follows from by the same argument used for $f$ convex in Lemma 2.4 by replacing ${\mathbb{R}}^{n}$ with $Y^{\varepsilon}$, which completes part (a).
For property (b), note that for each $y_{1},y_{2}$ in $Y^{\varepsilon}$ and each $\varepsilon$ small we have
$|u_{\varepsilon}(y_{1})-u_{\varepsilon}(y_{2})|\leqslant\int_{{\mathbb{R}}^{n}}K|y_{1}-y_{2}|\,\eta_{\varepsilon}(z)\,\mathrm{d}z=K|y_{1}-y_{2}|,$
where we have used that $u$ is $K$-Lipschitz and $\eta$ is non-negative with integral equal to $1$.
It remains only to show the decreasing limit property of part (c), which is a consequence the convexity of $u$ and the properties of the mollifier $\eta$. First note that, since $u$ is convex and $\eta$ is even (that is, (10.1) holds),
$\begin{split}u(y)=\int_{{\mathbb{R}}^{n}}u(y)\eta_{\varepsilon}(z)\,\mathrm{d}z&\leqslant\frac{1}{2}\int_{{\mathbb{R}}^{n}}u(y-z)\eta_{\varepsilon}(z)\,\mathrm{d}z+\frac{1}{2}\int_{{\mathbb{R}}^{n}}u(y+z)\eta_{\varepsilon}(z)\,\mathrm{d}z\\ &\leqslant\frac{1}{2}\int_{{\mathbb{R}}^{n}}u(y-z)\eta_{\varepsilon}(z)\,\mathrm{d}z+\frac{1}{2}\int_{{\mathbb{R}}^{n}}u(y-z)\eta_{\varepsilon}(-z)\,\mathrm{d}z=u_{\varepsilon}(y),\end{split}$
and hence $u\leqslant u_{\varepsilon}$ on $Y^{\varepsilon}$. Given the pointwise (locally uniform) convergence of $u_{\varepsilon}$ to $u$, it remains only to show the monotonicity
$0<\delta<\varepsilon\quad\implies\quad u_{\delta}(y)\leqslant u_{\varepsilon}(y)\quad\forall\,y\in Y^{\varepsilon}\subset Y^{\delta}.$ (10.4)
The claim (10.4) is a simple consequence of the following four-point inequality for the convex function $u$: if $0<\delta<\varepsilon$ then
$u(y-\delta z)+u(y+\delta z)\leqslant u(y-\varepsilon z)+u(y+\varepsilon z),$ (10.5)
which holds for each $y\in Y^{\varepsilon}$ if $\varepsilon$ is sufficiently small. To prove (10.5), it is enough to note that with
$t=\frac{\varepsilon+\delta}{2\varepsilon}\in(0,1)\quad\text{and}\quad 1-t=\frac{\varepsilon-\delta}{2\varepsilon}\in(0,1)$
one has
$y-\delta z=t(y-\varepsilon z)+(1-t)(y+\varepsilon z)\quad\text{and}\quad y+\delta z=t(y+\varepsilon z)+(1-t)(y-\varepsilon z).$
For the decreasing property (10.4), starting from (10.3), changing variables and using that $\eta$ is even, one finds that
$\begin{split}u_{\varepsilon}(y)&=\int_{B_{\varepsilon}(0)}u(y-z)\eta\!\left(\frac{z}{\varepsilon}\right)\varepsilon^{-n}\,\mathrm{d}z=\int_{B_{1}(0)}u(y-\varepsilon z)\eta(z)\,\mathrm{d}z\\ &=\int_{B_{1}(0)\cap\{z_{1}>0\}}\bigl{(}u(y-\varepsilon z)+u(y+\varepsilon z)\bigr{)}\,\eta(z)\,\mathrm{d}z;\end{split}$
using this formula also for $u_{\delta}$, the four-point inequality (10.5) gives (10.4). ∎
We proceed now with the simplifications. First, note that the result of Theorem 10.1 is local. If one were to use directly the Definition 2.1, the proof would entail showing that for each pair of points $(x,y)\in X\times Y$ with $y=\Phi(x)$ and $u$ being $\lambda$-semiconvex on some ball $B_{\rho}(y)\Subset Y$ for some $\lambda>0$, one has $v\vcentcolon=u\circ\Phi$ is $\Lambda$-semiconvex on some ball $B_{\sigma}(x)\subset\Phi^{-1}(B_{\rho}(y))\Subset X$ for some $\Lambda>0$. Instead, we exploit this local nature to make the first two simplifying assumptions in the statement of the theorem. By replacing $Y$ with some smaller ball than $B_{\rho}(y)$, we can assume that
@@
where $\mathcal{N}(\overline{X})$ and $\mathcal{N}(\overline{X})$ are open neighborhoods of $\overline{X}$ and $\overline{Y}$ both compact and
$u$ is $\lambda$-semiconvex on $\mathcal{N}(\overline{Y})$ convex. (10.7)
With these reductions, it is enough to prove that
$v=u\circ\Phi$ is $\Lambda$-semiconvex on $X$, for some $\Lambda>0$. (10.8)
We can further simplify by assuming that
$u$ is convex on $\mathcal{N}(\overline{Y})$ convex. (10.9)
To see this, replace $u$ by $u_{\lambda}(y)\vcentcolon=u(y)+\frac{\lambda}{2}|y|^{2}$, which is convex on $\overline{Y}$ by (10.7). Since the pullback function
$v_{\lambda}(x)\vcentcolon=u_{\lambda}(\Phi(x))=u(\Phi(x))+\frac{\lambda}{2}|\Phi(x)|^{2}=v(x)+\frac{\lambda}{2}|\Phi(x)|^{2},$
differs from $v$ by the smooth function $\varphi(x)\vcentcolon=\frac{\lambda}{2}|\Phi(x)|^{2}$, we can make use of the standard fact that if $\varphi\in C^{2}(X)$
$v$ is locally-semiconvex on $X\ \ \Leftrightarrow\ \ v+\varphi$ is locally-semiconvex on $X$,
which follows from the local boundedness of the Hessian of $\varphi\in C^{2}(X)$.
Next, since $u$ is convex on the convex neighborhood $\mathcal{N}(\overline{Y})$, by replacing $\mathcal{N}(\overline{Y})$ with a smaller convex neighborhood of $\overline{Y}$ compact, we can assume that
@@
which follows from Theorem 1.16(b).
With these reductions, it suffices to show (10.8), assuming (10.6), (10.9) and (10.10); that is, assuming that $\Phi:\mathcal{N}(\overline{X})\to\mathcal{N}(\overline{Y})$ is a diffeomorphism and
[…]
We are now ready for the last reduction, in which we can also assume that
[…]
To do that, consider the potential theory determined by the following subequation:
=$\mathcal{F}=\{(r,p,A)\in{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n):\ |p|\leqslant K,\ A\geqslant 0\}.$
[…]
Considering the simplifications provided by (10.11) and (10.12), proving Theorem 10.1 now amounts to showing that the following result holds.
Lemma 10.3.
Suppose that $\Phi:\mathcal{N}(\overline{X})\to\mathcal{N}(\overline{Y})$ is a diffeomorphism between open neighborhoods of $\overline{X}$ and $\overline{Y}$, where $X$ and $Y$ are open and relatively compact in ${\mathbb{R}}^{n}$. Assume (10.11) and (10.12); that is,
$u$ is smooth, convex and $K$-Lipschitz on $\mathcal{N}(\overline{Y})$ convex.
Define the constant [^1] Note that $\mathcal{N}(\overline{X})\ni x\mapsto\langle D^{2}\Phi(x)e,e\rangle=\sum_{i,j=1}^{n}\frac{\partial^{2}\Phi(x)}{\partial x_{i}\partial x_{j}}\,e_{i}e_{j}\in{\mathbb{R}}^{n},$ being $\Phi=(\Phi_{1},\dots,\Phi_{n})$.
$C\vcentcolon=\sup\bigl{\{}\bigl{|}\langle D^{2}\Phi(x)e,e\rangle\bigr{|}:\ x\in\overline{X},\ |e|=1\bigr{\}}.$ (10.13)
Then $v\vcentcolon=u\circ\Phi$ is $\Lambda$-semiconvex on $X$ for $\Lambda=KC$.
Proof.
Since $v$ is smooth, it is enough to show that $D^{2}v(x)+\Lambda I\geqslant 0$ on $X$, which holds if we can show that
$\langle D^{2}v(x)e,e\rangle\geqslant-KC\quad\forall x\in\overline{X},\ \forall e\in\mathbb{S}^{n-1}.$
By applying twice the chain rule to $v=u\circ\Phi$ one finds
$\begin{split}\langle D^{2}v(x)e,e\rangle&=\sum_{i,j=1}^{n}\frac{\partial^{2}v}{\partial x_{i}\partial x_{j}}(x)e_{i}e_{j}\\ &=\langle(D^{2}u(\Phi(x))(D\Phi(x)e),D\Phi(x)e\rangle+\sum_{k=1}^{n}\frac{\partial u}{\partial y_{k}}(\Phi(x))\langle D^{2}\Phi_{k}(x)e,e\rangle.\end{split}$
Since $u$ is convex, the first term is non-negative. It remains to show that the second term is bounded below by $-KC$. This follows from the Cauchy–Schwartz inequality, since $u$ is $K$-Lipschitz and recalling the definition of $C$ (cf. (10.13)). ∎
Hence, Theorem 10.1 is also proven. As a consequence of it, one can give the following definition.
Definition 10.4.
Let $M$ be an $n$-dimensional differentiable manifold, and let $u\colon M\to{\mathbb{R}}$. We say that $u$ is locally semiconvex on $M$ if $u\circ\varphi^{-1}$ is locally semiconvex on $\varphi(U)\subset{\mathbb{R}}^{n}$ for any chart $(U,\varphi)$ of $M$.
Notice that Theorem 10.1 guarantees that such a definition is chart-independent; indeed, given two charts $(U,\varphi)$ and $(V,\psi)$ with $U\cap V\neq\emptyset$, then $u\circ\varphi^{-1}|_{\varphi(U\cap V)}$ is locally semiconvex if and only if $u\circ\psi^{-1}|_{\psi(U\cap V)}$ is locally semiconvex, since $\Phi\vcentcolon=\varphi\circ\psi^{-1}|_{\psi(U\cap V)}\colon\psi(U\cap V)\to\varphi(U\cap V)$ is a diffeomorphism.
Appendix A The Legendre transform and Alexandrov’s theorem
This appendix, together with the next appendix, provides a self-contained proof of Alexandrov’s Theorem 1.29 on second order differentiability of convex functions. The proof is based on two ingredients. The first ingredient concerns properties of the Legendre transform, which is well known to be an important tool in the study of convex functions. Roughly speaking, this transform provides a means of passing from first-order differentiability to second-order differentiability for convex functions. The second ingredient is a Lipschitz version of Sard’s theorem (Theorem A.10).
Our proof of Alexandrov’s theorem follows the main steps of Harvey and Lawson’s proof in [28], which are basically the same as those proposed by Crandall, Ishii, and Lions in [14]. A different proof exploiting Lebesgue’s Differentiation Theorem and mollifications is given in the monograph of Evans and Gariepy [19].
We begin with the definition of Legendre transform. Given $f\colon X\to{\mathbb{R}}$, with $X\subset{\mathbb{R}}^{n}$, recall the Definition 1.7 of its subdifferential $\partial f$.
Definition A.1.
The Legendre transform of $f$ is the function $g\colon\partial f(X)\to{\mathbb{R}}$ defined by the relation
$f(x)+g(y)=\langle x,y\rangle\qquad\forall(x,y)\in\partial f;$ (A.1)
that is, one defines
$g(y)\vcentcolon=\langle x,y\rangle-f(x)\quad\forall\,y\in\partial f(x),\ x\in X$ (A.2)
We will also use the notation $g=\mathscr{L}f$.
Remark A.2 (Alternate definition of the Legendre transform).
@@ Appendix A The Legendre transform and Alexandrov’s theorem
Notice that the definition (A.2) is somewhat implicit; that is, in order to define $g(y)=\mathscr{L}f(y)$ for a given $y\in\partial f(X)={\rm dom}(g)$, one must identify a fiber $\partial f(x)$ in which $y$ lives. A more explicit way to define the Legendre transform $g$ of $f$ involves an optimization procedure based on the observation that
$(x,y)\in\partial f\quad\iff\quad f(z)-\langle y,z\rangle\geqslant f(x)-\langle y,x\rangle\quad\forall z\in X;$
@@
We will see that these equivalent formulations (A.2) and (A.3) are both useful.
Remark A.3.
If we consider (,+]($\infty$,$\infty$]( - ∞ , + ∞ ]-valued functions (which are not identically equal to +$\infty$+ ∞), then the identity in (A.3) provides a definition of $g=\mathscr{L}f$ for all $y\in{\mathbb{R}}^{n}$, with possibly $g(y)=+\infty$ for $y\notin\partial f(X)$. This corresponds to the convex conjugate of the extension $\tilde{f}$ of $f$ defined by letting $\tilde{f}=+\infty$ outside $X$.
@@
Let $f:X\to{\mathbb{R}}$, with $X\subset{\mathbb{R}}^{n}$, and suppose that $\partial f(X)\neq\emptyset$. Then the following hold:
@@
    $\mathscr{L}f(y)=\sup_{z\in X}\left(\langle y,z\rangle-f(z)\right)<+\infty;$ (A.4)
@@
    for each $x\in X$ such that $\partial f(x)\neq\emptyset$ and $y\in\partial f(X)$,
    $(x,y)\in\partial f\quad\iff\quad(y,x)\in\partial(\mathscr{L}f);$
@@
    $\tilde{X}\vcentcolon=\{x\in X:\,\partial f(x)\neq\emptyset\}\subset\partial(\mathscr{L}f)(\partial f(X))$ and ${\left.\kern-1.2pt\mathscr{L}(\mathscr{L}f)\vphantom{|}\right|_{\tilde{X}}}=f$.
@@
The equivalence of part (a) has been presented in Remark A.2. Part (b) follows from the fact that $\mathscr{L}f$ is the supremum of the family of affine functions $\{\langle\cdot,z\rangle-f(z)\}_{z\in X}$; indeed, the convexity of $\mathscr{L}f$ comes from the property (c) recalled in Remark 1.6 (also note that $\partial f(X)$ is convex by Lemma 1.12(a), since $\partial f(x)$ is convex even when $f$ is not, as noted in the proof).
For the implication ()($\Longrightarrow$)( ⟹ ) of part (c), let $(x,y)\in X\times\partial f(X)$. Suppose that $(x,y)\in\partial f$; that is, $y\in\partial f(x)$. In order to show that $(y,x)\in\partial g$, where $g=\mathscr{L}f$, one needs to show that $y\in\partial f(x)$; that is,
@@
$f(z)+\langle w,\hat{x}-z\rangle\leqslant f(\hat{x}),\quad\forall\,\hat{x}\in X;$ (A.6)
by the definition (A.2) of the Legendre transform we have $g(w)\vcentcolon=\langle z,w\rangle-f(z)$ and $g(y)\vcentcolon=\langle x,y\rangle-f(x)$, which imply
@@
where the last inequality uses (A.6) in $\hat{x}=x$, thus giving (A.5). Notice that we also have $\tilde{X}\vcentcolon=\{x\in X:\,\partial f(x)\neq\emptyset\}\subset\partial g(\partial f(X))$.
For the involution claim of part (d), notice that we have shown that $(x,y)\in\partial f$ implies that $(y,x)\in\partial g$, where $g$ is convex. Hence, for each $x\in\tilde{X}$, by the definition (A.1) we have
$\mathscr{L}(\mathscr{L}f)(x)=\mathscr{L}g(x)\vcentcolon=\langle x,y\rangle-g(y)\vcentcolon=\langle x,y\rangle-\langle y,x\rangle+f(x)=f(x),$
which proves (d).
@@
$\langle x,y\rangle=g(y)+\mathscr{L}g(x)=\mathscr{L}f(y)+f(x),$
@@
$\langle x,y\rangle=\sup_{z\in X}\bigl{(}\langle y,z\rangle-f(z)\bigr{)}+f(x),$
yielding $f(z)\geqslant f(x)+\langle y,z-x\rangle$ for each $z\in X$; that is, $(x,y)\in\partial f$. ∎
@@
The properties (c) and (d) of Proposition A.4 simplify when $f$ is convex and $X$ is open, in which case $\tilde{X}=X$. However, it remains false, in general, that
$(x,y)\in\partial f\ \iff\ (y,x)\in\partial g,$
@@
It can in fact happen that $\partial g(\partial f(X))\supsetneqq X$, which also implies that $\mathscr{L}\!\mathscr{L}f\neq f$, in the sense that the domain of the former is larger than that of the latter. For example, consider $f=|\cdot|$, defined on $X=(-1,1)\subset{\mathbb{R}}$; it is easy to see that
$\partial f(x)=\begin{cases}-1&\text{if $x\in(-1,0)$}\\ [-1,1]&\text{if $x=0$}\\ 1&\text{if $x\in(0,1)$},\end{cases}$
so that, according to (A.1), $g\vcentcolon=\mathscr{L}f\equiv 0$ on $\partial f(X)=[-1,1]$. Therefore,